OpenStudy (anonymous):

Help On This One Please (: Derive the equation of the parabola with a focus at (0, 1) and a directrix of y = −1. f(x) = −one fourth x2 f(x) = one fourth x2 f(x) = −4x2 f(x) = 4x2

OpenStudy (anonymous):

@johnweldon1993 any idea how to do this lol :/ sorry im bugging

OpenStudy (johnweldon1993):

Sure! And dont worry about bugging me :) "focus of the parabola at (0, 1) directrix = -1 The focus lies above the directrix, so the parabola opens upwards. ...and we know that the equation of an UP parabola is y = a(x - h)² + k The vertex is halfway between focus and directrix so vertex (0, 1) h = 0 k = 1 p = distance between focus and vertex = 2 a = 1/(4p) = 1/4 The equation of the parabola: y = (1/4)(x^2) + 1 y = (1/4)x^2 So it would be the one forth x^2